How do you solve two problems at the same time and get a coordinationes
example 2c-d=-2 second problem is 4c+d=20

Answer by stanbon(75887)

2c-d=-2
4c+d=20
--------
This is called a system of equations.
You are asked to find a solution that
satisfies both equations.
---------
Add the two equation to solve for "c":
6c = 18
c = 3
----------------
Substitute into 2c - d = -2 to solve for "d":
2*3 - d = -2
d = 8
-------------------
Check the answer in both equations:
2c-d=-2
2*3 - 8 = -2
6 - 8 = -2
-2 = -2
--------------------
4c+d=20
4*3 + 8 = 20
12 + 8 = 20
20 = 20
====================
Both check out so the solution is c=3, d=8
============================================
Cheers,
Stan H.